OpenStudy (anonymous):

Help please!! Simplify in terms of sin(theta) and cos(theta) (1-cos(theta))(1+cos(theta))

OpenStudy (anonymous):

\[(1-\cos \theta)(1+\cos \theta)\]

OpenStudy (mertsj):

Multiply those two binomials.